Will is 8 years old and Jan is 22 in how many years will Jan be twice as old as will?
Answer by ankor@dixie-net.com(22740)

Will is 8 years old and Jan is 22
in how many years will Jan be twice as old as will?
:
let y = no. of yrs for this to be true
y + 22 = 2(y + 8)
y + 22 = 2y + 16
22 - 16 - 2y - y
6 = y
y = 6 yrs
:
:
Check this; in six yrs W will be 14 and J will be 28